OpenStudy (mimi_x3):

\[\int\frac{1}{ln(x+1)} dx\]

OpenStudy (mimi_x3):

@wasiqss

OpenStudy (wasiqss):

lemme try

OpenStudy (mimi_x3):

Hint: similar to your other one.

OpenStudy (mimi_x3):

But this is apparently an integral which does not have a close form..

OpenStudy (wasiqss):

okay

OpenStudy (wasiqss):

ash dont solve its for me

OpenStudy (ash2326):

I ain't solving, just for reference http://www.wolframalpha.com/input/?i=integral+1%2F%28ln+%28x%2B1%29%29+dx

OpenStudy (mimi_x3):

this is like wasiqs previous integral which was apparently solveable http://www.wolframalpha.com/input/?i=integral+e^x%2F%28x%29

OpenStudy (wasiqss):

mishoo it was easy ~~~ i did it

OpenStudy (wasiqss):

just apply by parts

OpenStudy (mimi_x3):

by parts? please elaborate.

OpenStudy (mimi_x3):

i thought the first step is a substitution..

OpenStudy (wasiqss):

yeh take 1 as first functio

OpenStudy (wasiqss):

\[x/{(\ln(x+1)} - \int\limits_{}^{}x^2 +x\]

OpenStudy (mimi_x3):

how about \(u= ln(x+1)\) => \(e^{u}= x+1\)

OpenStudy (wasiqss):

if it can be done easily why take substitution..

OpenStudy (mimi_x3):

but this is an integral which does not have a closed form right?

OpenStudy (wasiqss):

i dont understand forms :P i just know how to solve the ugliest of integrals :P

OpenStudy (mimi_x3):

lol, but look at the answer on wolfram..

OpenStudy (wasiqss):

i dont trust wolfram cause in two questions i pointed out its mistake and did what it cant do :P

OpenStudy (mimi_x3):

lol are you sure that you're right though? :P

OpenStudy (wasiqss):

i can do better than wolfram :P

OpenStudy (mimi_x3):

lol